# A Cloud in the Sky: Geo-Aware On-board Data Services for LEO Satellites

### 10 Oct 2024 v1

@@ 1 Introduction
With more connected devices and improved capabilities of connected devices, the need for high-bandwidth connectivity anywhere and anytime keeps growing at an exponential rate. High-population areas usually see investments in high-capacity communication infrastructure such as cell towers, small-cell antennas, radio heads and base stations. For an end-user device wanting access to a wireless mobile network, having low-latency access to this infrastructure not only allows the initial connection to be established faster, but also encounters fewer restrictions on the effective throughput, given timeouts of various acknowledgments in the protocols.
@@
We propose to host a bandwidth ledger that enables on-demand purchases of cellular bandwidth akin to the way compute resources are purchased in the Cloud, and that allows revenue sharing between the MNOs and SNOs. Given the latency concerns described above, we further propose to host this ledger entirely on-board the LEO satellites, using their Inter-Satellite-Link (ISL) free-space optical communication links.
@@ 3 Related Work
The many challenges of offering cellular connectivity from satellites, including doppler effects and latency issues, are described in [8]. These and other challenges in using the traditional satellite communication spectrum in the L-Band and S-Band for direct-to-cell connectivity [17] are the reason for SNOs and MNOs to coordinate their usage and allocation of spectrum resources. FCC has also recognized the value of re-using MNO spectrum for extended or supplemental coverage through satellites [^1] https://www.fcc.gov/document/fcc-proposes-framework-facilitate-supplemental-coverage-space-0.
@@
We have previously envisioned a similar architecture for ad-hoc multi-provider bandwidth purchases using eSIMs and an open distributed ledger market in [25] and [23] to overcome the inefficiencies of establishing new roaming contracts between providers. The advantage of our approach compared to that of [13] is that contracts do not need to be negotiated ahead of time and there is no need to keep tokens on SIM cards. However, our approach requires the presence of a distributed ledger accessible to both the end-user device and the mobile operator, and thus the performance of the system depends on the placement and migration of the distributed ledger.
@@ 4 Blockchains and Distributed Ledgers
Blockchain technology was popularized with the virtual bitcoin currency [15] and provides a way to maintain a distributed ledger consistently and scalably across a large number of nodes. The 2-phase-commit transaction protocols used in traditional relational databases [2], for instance, scale very poorly as the transaction as a whole fails if only one party maintaining state fails. Blockchains originally provided a way to ensure consensus by what is referred to Proof of Work (PoW) where only parties that solved complex math problems were allowed to write into the ledger. The way a blockchain serializes blocks and maintains a hash of the previous block in the hash of the current block, it is easy to validate the internal consistency of the chain. PoW blockchains may contain competing forks of the blockchain but the longest chain wins if there is a conflict. These types of blockchains are appropriate in environments where anyone is allowed to write into the ledger in a large community of untrusted parties. The computation to solve math problems is however energy hungry and since LEO satellites have a limited lifetime depending on how long their battery lasts, this type of blockchain, referred to as permissionless, is not appropriate for our use case.
Instead we focus on permissioned blockchains that only allow authenticated and trusted participants to write into the ledger. Hence the blockchain can never fork into competing branches and no work is wasted. Some nodes may have fewer blocks but they will eventually catch up. Still the protocol needs to deal with failures and needs to ensure a consistent ordering of transactions. This is typically done by having a single node act as the leader to provide consistent ordering. Electing a leader is a critical part of the protocol as having no or multiple leaders will cause the system to fail as a whole. Leader election can be done by various consensus algorithms such as Paxos [12] or Raft [16]. As we shall see later we can provide a more suited and simpler leader election algorithm for the LEO case that is similar to the token ring leader election algorithm [27].
@@
This process allows writes and reads to be done with eventual consistency [30], or optimistic locking as opposed to with strict atomicity, consistency, isolation and durability (ACID) guarantees, known from protocols such as two-phase commit [2]. As previously mentioned, the issue with the ACID protocols is that they scale poorly across large sets of replicas as failures become more likely.
@@ 5 LEO Constellations
Each satellite would typically have 4 inter-satellite links to neighboring satellites, e.g. west, east, north and south, using free-space optical connections to transmit at the speed of light. The network in a shell can thus be seen as a connected mesh forming a 2d-torus where there is a wrap-around both in rows and columns (see Figure 1). This architecture is referred to as a +GRID inter-satellite link (ISL) network [18], and it is the one we assume for our work.
@@
For example, the Pacific Ocean stretches about 15,50015,500km from the Arctic to the Southern Ocean and 19,80019,800km from Indonesia to the coast of Colombia, which is close to half the circumference of the earth. So up to a third of the planes and half of all the satellites in a plane could hover over the Pacific Ocean at any given time. A typical LEO constellation would make up about 22-28 orbital planes with 5-72 satellites per plane or 375-1584 satellites in total [^5] These numbers are based on the data in Table I in [18]. So about 8 planes and 39 satellites per plane or about 321 satellites could be wasting their capacity while idling over the Pacific. Having 300+ nodes in a distributed ledger with virtually unlimited speeds to interconnect them is clearly a resource worth exploiting.
One major issue with these networks is that adding more capacity by sending up more satellites to increase coverage also means that there is more waste in terms of periods where the satellites are idle. Furthermore, these periods are predictable and cyclical so easy to identify and exploit without any central coordination666each satellite node can determine independently when they are above a certain geographic area without a need for external communication, which is the key idea behind our approach, described next.
@@ 6 Protocol
All our protocols rely on the 2d-torus architecture of LEO ISL. Furthermore we define a blockchain Service Area (SA) as a block of contiguous nodes (each node being hosted on a satellite) over some geographic area with low terrestrial traffic load, such as the Pacific Ocean. All satellites in the constellation may communicate with each other through the one-hop architecture of the 4 ISL links (see Figure 1), but broadcasting is only done in the service area, and only nodes in the service area take active part in processing transactions and maintaining up-to-date blockchain state. Owing to the orbital movement, the set of nodes that are in the service area keeps changing, so we need to do continuous migrations of the blockchain ledger to nodes that move into the service area. Note that the nodes currently in the service area do not need to be informed about the new nodes moving into the service area, because the movement of every satellite is completely predictable and so each node in the service area can independently determine when it will move out of the service area and which node will enter to take its place.
@@ 6.1 Gossip and Routing
Updates on one node need to be broadcast efficiently to other nodes in the blockchain cluster to ensure the window of inconsistency is kept small. At the same time a single path to distribute updates can be a single point of failure, e.g. if all updates go through a central node. So, for reliability most blockchain systems implement some form of gossip protocol [7]. These protocols are modeled after epidemic virus infection spread in a human population: if infected, you then infect your nearest (uninfected) neighbors who go on to infect their (uninfected) neighbors, while an encounter between two infected people does not propagate the virus further. Typically a handful of neighbor nodes are configured to propagate gossip messages to. Care is taken not to propagate to a neighbor that sent you the broadcast message. A message that has already been seen, e.g. previously received by a different neighbor, is not propagated again.
The gossip protocol is also efficient, in the sense that $N$ nodes can be synchronized in $O(\log N)$ iterations despite node failures and data losses in transmission [7]. The gossip protocol on a torus (which is the scenario that matches LEO satellite constellations) was analyzed in [14]. Improved gossip protocols specifically for use in blockchains were proposed and evaluated in [1][11], and [21].
In a LEO constellation architecture following the +GRID ISL structure the neighbors are simply the north, south, east, and west connections. If a message comes from one direction it is propagated into the other three, and previously seen messages are dropped. Furthermore we restrict propagation to within the service area we defined, typically a grid but it could have any shape as long as all the nodes in the service area are contiguous. All messages traversed also receive the gossip due to the one-hop communication structure, ensuring the entire service area receives a broadcast efficiently.
Communication within an orbital plane tends to be more reliable and thus more nodes are reached within a time unit if west/east spreading is prioritized but propagation can be done concurrently in all directions.
@@
For instance a transaction may be created from any node but needs to be executed in a node in the current service area and the leader node that orders transactions then broadcasts the ordered transactions within the service area only.
@@ 6.2 Transaction Processing
Transaction processing is at the core of our protocol as it is what ensures eventually consistent distributed states and availability even in the case of partial failures. A survey of blockchain consensus algorithms may be found in [9]. Distributing state across a large area becomes even more important in an ISL network where only single node hops are possible. As we have previously mentioned it is enough to contact any node (satellite) currently in the service area to read the current state.
@@
    If local execution succeeded, the node forwards the transaction to the leader node for ordering.
@@
    When submitting a transaction, a transaction ID is generated and the submitter may query any node in the service area to check whether a transaction completed successfully. A successfully completed transaction including all its read and write operations are written into the blockchain and the state is updated accordingly.
@@ 6.3 Service Area Migration
    Finally when the migration is complete the west-most nodes in the service area orbits can drop out, and the borders of the service area can be moved to include the nodes that just migrated into it. Note that the timing of these steps can be predicted within each node, based on the location of the nodes, and thus there is no need to send messages to trigger these steps.
@@ 6.4 Smart Contracts
To be able to test our transaction processing infrastructure better, we also provide a smart contract programming construct that allows us to define smart contracts. A smart contract is simply an interface with methods that read and write from the state maintained in the blockchain, as well as define what the structure of the state is. Each method when executed generates a transaction comprising an ordered list of read and write operations that can be submitted into the constellation for processing and that will be written to the ledger with the eventually consistent guarantees.
@@ 7 Simulation and Visualization
To be able to test transaction processing and smart contract execution in a constellation in motion we developed a simulation using Python and a Web visualization.

### 14 Oct 2024 v2

@@ 1 Introduction
With so many connected devices and their improved capabilities, the need for high-bandwidth connectivity anywhere and anytime keeps growing at an exponential rate. As a result, high-population areas usually see investments in high-capacity communication infrastructure such as cell towers, small-cell antennas, radio heads and base stations. For an end-user device wanting access to a wireless mobile network, having low-latency access to this infrastructure not only allows the initial connection to be established faster, but also encounters fewer restrictions on the effective throughput, given timeouts of various acknowledgments in the protocols.
@@
In this paper, we propose to host a bandwidth ledger that enables on-demand purchases of cellular bandwidth akin to the way compute resources are purchased in the Cloud, while allowing revenue sharing between the MNOs and SNOs. Given the latency concerns described above, we further propose to host this ledger entirely on-board the LEO satellites, using their Inter-Satellite-Link (ISL) free-space optical communication links.
@@ 3 Related Work
The many challenges of offering cellular connectivity from satellites, including doppler effects and latency issues, are described in [8]. These and other challenges in using the traditional satellite communication spectrum in the L-Band and S-Band for direct-to-cell connectivity [17] are the reason for SNOs and MNOs to coordinate their usage and allocation of spectrum resources.
The FCC has also recognized the value of re-using MNO spectrum for extended or supplemental coverage through satellites [^1] https://www.fcc.gov/document/fcc-proposes-framework-facilitate-supplemental-coverage-space-0.
@@
We have previously envisioned a similar architecture for ad-hoc multi-provider bandwidth purchases using eSIMs and an open distributed ledger market in [25] and [23] to overcome the inefficiencies of establishing new roaming contracts between providers. The advantage of our approach compared to that of [13] is that contracts do not need to be negotiated ahead of time and there is no need to keep tokens on SIM cards. Instead, our approach requires the presence of a distributed ledger accessible to both the end-user device and the mobile operator, and thus the performance of the system depends on the placement and migration of the distributed ledger.
@@ 4 Blockchains and Distributed Ledgers
Blockchain technology was popularized with the virtual bitcoin currency [15] and provides a way to maintain a distributed ledger consistently and scalably across a large number of nodes. The 2-phase-commit transaction protocols used in traditional relational databases [2], for instance, scale very poorly as the transaction as a whole fails if only one party maintaining state fails. Blockchains originally provided a way to ensure consensus by what is referred to Proof of Work (PoW) where only parties that solved complex math problems were allowed to write into the ledger. The way a blockchain serializes blocks and maintains a hash of the previous block in the hash of the current block, it is easy to validate the internal consistency of the chain. PoW blockchains may contain competing forks of the blockchain but the longest chain wins if there is a conflict. These types of blockchains are appropriate in environments where anyone is allowed to write into the ledger when partaking in a large community of untrusted parties. The computation to solve math problems is however energy hungry and since LEO satellites have a limited lifetime depending on how long their battery lasts, this type of blockchain, referred to as permissionless, is not appropriate for our use case.
Instead we focus on permissioned blockchains that only allow authenticated and trusted participants to write into the ledger. Hence the blockchain can never fork into competing branches and no work is wasted. Some nodes may have fewer blocks but will eventually catch up. Still, the protocol needs to deal with failures while ensuring a consistent ordering of transactions. This is typically done by having a single node act as the leader to provide consistent ordering. Electing a leader is a critical part of the protocol, as having no leader or multiple ones will cause the system to fail as a whole. Leader election can be done by various consensus algorithms such as Paxos [12] or Raft [16]. As we shall see later we can provide a more suited and simpler leader election algorithm for the LEO case that is similar to the token ring leader election algorithm [27].
@@
This process allows reads and writes to be done with eventual consistency [30], or optimistic locking as opposed to with strict atomicity, consistency, isolation and durability (ACID) guarantees, known from protocols such as two-phase commit [2]. As previously mentioned, the issue with the ACID protocols is that they scale poorly across large sets of replicas as failures become more likely.
@@ 5 LEO Constellations
Each satellite typically has 4 inter-satellite links to neighboring satellites, e.g. west, east, north and south, using free-space optical connections to transmit at the speed of light. The network in a shell can thus be seen as a connected mesh forming a 2d-torus where there is a wrap-around both in rows and columns (see Figure 1). This architecture is referred to as a +GRID inter-satellite link (ISL) network [18], and it is the one we assume for our work.
@@
For example, the Pacific Ocean stretches about 15,50015,500km from the Arctic to the Southern Ocean and 19,80019,800km from Indonesia to the coast of Colombia, which is close to half the circumference of the earth. So up to a third of the planes and half of all the satellites in a plane could hover over the Pacific Ocean at any given time. A typical LEO constellation would make up about 22-28 orbital planes with 5-72 satellites per plane or 375-1584 satellites in total [^5] These numbers are based on the data in Table I in [18]. Thus about 8 planes and 39 satellites per plane or about 321 satellites could be wasting their capacity while idling over the Pacific. Having 300+ nodes in a distributed ledger with virtually unlimited speeds to interconnect them is clearly a resource worth exploiting.
One major issue with these networks is that adding more capacity by sending up more satellites to increase coverage also means that there is more waste in terms of periods where the satellites are idle. Furthermore, these periods are predictable and cyclical thus easy to identify and exploit without any central coordination666each satellite node can determine independently when they are above a certain geographic area without a need for external communication, which is the key idea behind our approach, which we describe next.
@@ 6 Protocol
All our protocols rely on the 2d-torus architecture of LEO ISL. Furthermore we define a blockchain Service Area (SA) as a block of contiguous nodes (each node being hosted on a satellite) over some geographic area with low terrestrial traffic load, such as the Pacific Ocean. All satellites in the constellation may communicate with each other through the one-hop architecture of the 4 ISL links (see Figure 1), but broadcasting is only done in the service area, and only nodes in the service area take active part in processing transactions and maintaining up-to-date blockchain state. Owing to their orbital movement, the set of nodes that are in the service area keeps changing, so we need to perform continuous migrations of the blockchain ledger to nodes that move into the service area. Note that the nodes currently in the service area do not need to be informed about the new nodes moving into the service area, because the movement of every satellite is completely predictable. Thus, each node in the service area can independently determine when it will move out of the service area and which node will enter to take its place.
@@ 6.1 Gossip and Routing
Updates on one node need to be broadcast efficiently to other nodes in the blockchain cluster to ensure the window of inconsistency is kept small. At the same time a single path to distribute updates can be a single point of failure, e.g. if all updates go through a central node. So, for reliability most blockchain systems implement some form of the gossip protocol [7]. These protocols are modeled after epidemic virus infection spread in a human population: if infected, you then infect your nearest (uninfected) neighbors who go on to infect their (uninfected) neighbors, while an encounter between two infected people does not propagate the virus further. Typically a handful of neighbor nodes are configured to propagate gossip messages to each other, while avoiding to propagate to a neighbor that sent you the broadcast message. A message that has already been seen, e.g. previously received by a different neighbor, is not propagated again.
The gossip protocol is also efficient, in the sense that $N$ nodes can be synchronized in $O(\log N)$ iterations despite node failures and data losses in transmission [7]. The gossip protocol on a torus (which is the scenario that matches LEO satellite constellations) was analyzed in [14]. Improved gossip protocols designed specifically for use in blockchains were proposed and evaluated in [1][11], and [21].
In a LEO constellation architecture that follows the +GRID ISL structure the neighbors are simply the north, south, east, and west connections. If a message comes from one direction it is propagated into the other three, and previously seen messages are dropped. Furthermore we restrict propagation to within the service area we defined, typically a grid but it could have any shape as long as all the nodes in the service area are contiguous. All messages traversed also receive the gossip due to the one-hop communication structure, ensuring the entire service area receives a broadcast efficiently.
Communication within an orbital plane tends to be more reliable and thus more nodes are reached within a time unit if west/east spreading is prioritized, but propagation can be done concurrently in all directions.
@@
For instance, a transaction may be created from any node but has to be executed in a node in the current service area. In that case the leader node that orders transactions then broadcasts the ordered transactions within the service area only.
@@ 6.2 Transaction Processing
Transaction processing is at the core of our protocol as it is what ensures eventually consistent distributed states and availability even in the case of partial failures. A survey of blockchain consensus algorithms may be found in [9]. Distributing state across a large area becomes even more important in an ISL network where only single node hops are possible. As we previously mentioned it is enough to contact any node (satellite) currently in the service area to read the current state.
@@
    If local execution succeeds, the node forwards the transaction to the leader node for ordering.
@@
    When submitting a transaction, a transaction ID is generated and the submitter may query any node in the service area to check whether a transaction completed successfully. A successfully completed transaction including all its read and write operations is then written into the blockchain and the state is updated accordingly.
@@ 6.3 Service Area Migration
    Finally, when the migration is complete the west-most nodes in the service area orbits can drop out, and the borders of the service area can be moved to include the nodes that just migrated into it. Note that the timing of these steps can be predicted within each node, based on the location of the nodes, and thus there is no need to send messages to trigger these steps.
@@ 6.4 Smart Contracts
In order to be able to test our transaction processing infrastructure better, we also provide a smart contract programming construct that allows us to define smart contracts. A smart contract is simply an interface with methods that read and write from the state maintained in the blockchain, as well as define what the structure of the state is. Each method when executed generates a transaction comprising an ordered list of read and write operations that can be submitted into the constellation for processing and that will be written to the ledger with the eventually consistent guarantees.
@@ 7 Simulation and Visualization
To be able to test transaction processing and smart contract execution in a constellation in motion, we developed a simulation using Python and a Web visualization.